# LLMR: Real-time Prompting of Interactive Worlds using Large Language Models

### 19 Dec 2023 v2

@@ 3.1. Planner
(1) $\begin{split}\mathcal{P}(x_{1},x_{2},...,x_{N}|u_{1},u_{2},...,u_{N},\Omega)&=\mathcal{P}(x_{1}|u_{1},u_{2},...,u_{N},\Omega)\prod_{n=1}^{N-1}\mathcal{P}(x_{n+1}|inlinehl\bm{x}^{n},u_{1},u_{2},...,u_{N},\Omega)\\ &=\mathcal{P}(x_{1}|u_{1},\Omega)\prod_{n=1}^{N-1}\mathcal{P}(x_{n+1}|\bm{x}^{n},u_{n+1},\Omega)\end{split}$
@@ 3.2. Scene Analyzer
To tackle these issues, we created a separate module termed the Scene Analyzer, which is a properly prompted LLM $\mathrm{A}(s|u,\Omega)$ that outputs a succinct summary of $\Omega$ conditional on the user request. On a high level, one can think of the Scene Analyzer as a means of perception that relays an abstraction of the environment for downstream processing. An illustration of the module is provided in Figure 4. Concretely, the output $s_{n}\sim\mathrm{A}(\cdot|u,\Omega_{n})$ is used to reparametrize the density at each sampling step:
@@ 3.3. Builder-Inspector
Algorithm 2 Generate Code With Inspection
1:$u$: user’s request, $s$: scene summary, $h$: additional hint. $\mathrm{B}(x|u,s,h):$ Builder; $\mathrm{I}(r,v|u,s,x)$: Inspector; T: maximum number of inspections.
2:$t\leftarrow 0$triangleright Initialization
3:$r_{0}\leftarrow\emptyset$
4:$v_{0}\leftarrow$ False
5:while $t<T$ and $v_{t}$ is false do
6:     $x_{t}\sim\mathrm{B}(\cdot|u,s,h,r_{t})$triangleright Builder writes code $x_{t}$
7:     $(r_{t},v_{t})\sim\mathrm{I}(\cdot|s,x)$triangleright Inspector checks code, outputs verdict $v_{t}$ and suggestion $r_{t}$
8:     $t\leftarrow t+1$
9:end while
10:return x
@@ 3.5. Skill Library
The creation of the Skill Library Module is motivated by two primary challenges. The first is the token size limitation imposed by the GPT architecture on the context, or ”metaprompt,” provided to the Builder. Typically, the Builder is presented with a comprehensive list of various APIs and plug-ins that could be employed to meet the user’s needs. As the range of available skills expands, this list lengthens, eventually surpassing GPT’s token size limit for public users. The second challenge lies in the Builder’s attention capacity, which appears to be limited. Even when we attempt to condense all the available skills into the Builder’s metaprompt, it struggles to keep track of a specific skill when the list becomes too lengthy. This limitation is further exacerbated by the necessity to include precise coding examples for each plug-in to ensure their effective utilization by GPT. To address these challenges, we created the Skill Library module, denoted as $\mathrm{L}(h|u)$, which serves as a centralized repository for all available skills and as an attention mechanism that retrieves only the skills relevant to a specific user prompt. We illustrate this module in Figure 6.
@@
As an illustrative example, consider a skill we created for GPT’s use, which leverages a combination of generative and contrastive models along with the Sketchfab API to source and integrate 3D models into a scene. While we delve into the specifics of this skill in the next section, it is worth noting that the Skill Library only receives a high-level summary of how this particular skill functions, along with similar descriptors for other skills. The actual examples needed to use this skill are then retrieved and supplied to the Builder for execution.
@@
This approach ensures that the Skill Library and the Builder work in tandem to efficiently and effectively generate code that fulfills the user’s request, while overcoming the token size and attention capacity limitations of LLMs.
@@ 4. Incorporating Existing Open-Source 3D Assets
The process of generating interactive 3D scenes often involves the creation and placement of various objects. For instance, a request to create an office space might be decomposed into the generation of a desk, chair, lamp, and clock. While it is possible to generate these objects using primitives, a method that works well even for composite objects like a car or an entire room (depicted in the car of Figure 8 and the kitchen of Figure 1), there is a need to leverage the intricate objects created by artists and 3D developers that exhibit high real-world fidelity. Previous work has utilized objects from Sketchfab (Roberts et al., 2022b, a) and used the priors of GPT to size them accordingly to the real world. However, this approach encounters challenges when the user prompts an object, say a clock, and Sketchfab offers 50 different clocks, only three of which are suitable for an office setting.
@@
1:$u$: user’s prompt; $o$: object in $u$; $T$: target image; $\bm{S}$: screenshots; $L(o,s_{i})$: language similarity; $V(T,s_{i})$: visual similarity.
2:$\bm{S}^{\prime}\leftarrow\text{Top 5 }s_{i}\in\bm{S}\text{ with highest }L(o,s_{i})$
3:$s^{*}\leftarrow\arg\max_{s_{i}\in\bm{S^{\prime}}}V(T,s_{i})$
4:return $s^{*}$
This process is repeated to generate entire scenes. Figure 7 illustrates this pipeline. There is potential for further exploration to improve this pipeline. For instance, selecting from the visual similarity space before the language similarity space might yield better results. Future work will involve human feedback to identify the workflow that maximizes the likeness between the 3D object loaded and the user’s intended object.
@@ 5. Memory Management
By default, language models generate new words based on all previously sampled tokens, a configuration that may not be ideal due to their finite context length. For instance, this may hinder the model’s ability to engage in extended conversations. To mitigate this, techniques such as dialogue summarization and distillation can be employed (Wang et al., 2023a; Bae et al., 2022; Huang et al., 2023). Additional research has delved into leveraging persistent memory and retrieving in-context examples from databases to enhance few-shot performance (Wang et al., 2023b; Zhong et al., 2023).
Module
@@
Scene Analzyer
@@
Memory Mode
Memory-less
Memory-less
Limited-memory
Memory-less
@@ 6.1. Installation
Our framework can be easily added to any existing Unity scenes. The framework consists of a unity package and a few additional open-sourced packages (such as GLTF loader and OpenAI), and the installation process takes only a few steps. This enables anyone with an OpenAI API key to try our framework. We are strong proponents of the adaptability of our framework, and so we are in the process of open-sourcing the foundational framework along with several examples on GitHub. Readers who wish to try our framework can try out the example playground scenes or can easily add our Unity package to their existing Unity projects. They would need to obtain an OpenAI API key, a copy of the Roslyn compiler and optionally an account for Sketchfab if they wish to automatically load existing assets. In the Appendix, we also provide the metaprompts used for each LLMR’s modules for transparency.
@@ 7.1. Game Design and Creativity
Figure 10. Accessible Interface Features in Action. A1 and A2 show how a user can prompt the system to adjust the color scheme of a kitchen scene for red-green color-blind compatibility. B1 and B2 demonstrate the activation of a magnifier tool which can be used to focus attention. C1 and C2 reveal the option to hide objects deemed not kid-friendly, enhancing safety and customization.
@@ 7.2. Accessibility and Adaptive Interface
Similar to the accessibility feature in 2D documents, our framework can also be prompted to make a 3D scene accessible and adaptive to different user needs and preferences. Figure 10 shows three examples of editing an existing virtual kitchen scene to different requests. For example, one can request to make the scene to be more friendly to red-green color-blind users. For someone who is near-sighted, they can prompt a magnifier tool that zooms into a particular part of the room. An architect can use our framework to figure out if the space is friendly for wheelchair users or make sure objects in the room are child-proof.
Figure 11. Spontaneous Creation of Teaching Guides A demonstration of creating a guide for operating a coffee machine in which LLMR animates a hand model to point out the various steps of the operation. Our framework allows for the rapid creation of such guides and furthermore allows users to ask questions that were not predicted by the instructor beforehand, with appropriate motions being animated on the fly.
Figure 12. Simulated Rescue Plan. HoloLens displays the automated generation of a simulation of a rescue plan using our framework. The guide shows an interactable 3D terrain, helicopter, and simulated wind, allowing rescue workers to visualize the flight path under different weather conditions.
@@ 7.3. Remote Assistance and Planning
In a remote training scenario, typically, creating such a training guide requires custom creation, from rigging a gesture to placing a UI element. An instructor can use our framework to automate the generation of a training guide from a list of instructions. (Figure 11). The trainee can then, for example, use an AR device that overlays information on the machine. As the trainee advances through the steps, they can ask questions directly to the guide where answers can be generated in the context of the trainee’s learning progress. In another scenario of remote rescue planning, helicopter operators can prompt a simulation of the flight path given several target locations and see how the flight path might be affected by different wind conditions (Figure 12).

### 22 Mar 2024 v3

@@ 3.1. Planner
(1) $\begin{split}&\mathcal{P}(x_{1},x_{2},...,x_{N}|u_{1},u_{2},...,u_{N},\Omega)=\mathcal{P}(x_{1}|u_{1},u_{2},...,u_{N},\Omega)\ \times\\ &\times\prod_{n=1}^{N-1}\mathcal{P}(x_{n+1}|\bm{x}^{n},u_{1},u_{2},...,u_{N},\Omega)\\ &=\mathcal{P}(x_{1}|u_{1},\Omega)\prod_{n=1}^{N-1}\mathcal{P}(x_{n+1}|\bm{x}^{n},u_{n+1},\Omega)\end{split}$
@@
Scene Analyzer module. This diagram illustrates the process flow of the Scene Analyzer module within a Large Language Model for Mixed Reality system. The flow begins with a virtual scene depicted in the bottom-left, showing a room with various objects such as a bear, a ball, and furniture. This visual scene is translated into a structured text format known as JSON, capturing the properties of each object like name, position, rotation, and scale. A user input, highlighted in a purple box with the text ”Make the bear twice as big”, feeds into the Scene Analyzer. The Analyzer processes this input along with the scene’s JSON representation to produce an output that identifies and prioritizes elements relevant to the user’s request. Here, the output is indicated by a goldenrod-colored box pointing to the term ”Relevant Object: ’Bear’” after being processed by GPT-4, suggesting that the Scene Analyzer has successfully isolated the bear object as the focus for further action by subsequent system modules such as the Builder. This step is crucial for optimizing the language model’s context window and ensuring focus on user-relevant scene elements.
@@ 3.2. Scene Analyzer
To tackle these issues, we created a separate module termed the Scene Analyzer, which is a properly prompted LLM $\mathrm{A}(s|u,\Omega)$ that outputs a succinct summary of $\Omega$ conditional on the user request. At a high level, one can think of the Scene Analyzer as a means of perception that relays an abstraction of the environment for downstream processing. An illustration of the module is provided in Figure 4. Concretely, the output $s_{n}\sim\mathrm{A}(\cdot|u,\Omega_{n})$ is used to reparametrize the density at each sampling step:
@@
Builder-Inspector paradigm. The image is a diagram of the ”Builder-Inspector loop” concept in a framework. The diagram is divided into several sections, including ”Input,” ”Builder,” and two iterations, ”Iteration 1” and ”Iteration 2,” with an ”Inspector” at the bottom. The Input section at the top left shows a UI interface with various buttons and a field to enter a command, labeled ”Skill: NA, Scene: Object ’Box’, Instruction: ’Make the box twice as big’”. The Builder section has an orange background and is connected to Iteration 1, which contains a code snippet with a comment ”// command: ’Make the box twice as big’”, and a line of code that incorrectly scales a ”Box” object. The Inspector, which has a purple background, evaluates the code from Iteration 1, providing a ”Verdict: FAIL” and a suggestion to use ”GameObject.Find()” instead. Iteration 2 shows a revised code snippet with the suggested correction. The Inspector then gives a ”Verdict: SUCCESS”. Arrows indicate the flow from Input to Builder, then to each iteration, and feedback from the Inspector. This diagram illustrates a feedback loop process where code is generated, inspected, and iterated upon until it passes inspection or reaches a maximum number of iterations, enhancing the quality of generated scripts.
@@ 3.3. Builder-Inspector
Input : $u$: user’s request, $s$: scene summary, $h$: additional hint.
Require : $\mathrm{B}(x|u,s,h)$: Builder;
$\mathrm{I}(r,v|u,s,x)$: Inspector;
T: maximum number of inspections.
$t\leftarrow 0$;
$r_{0}\leftarrow\emptyset$;
$v_{0}\leftarrow$ False;
while $t<T$ and $v_{t}$ is false do
       $x_{t}\sim\mathrm{B}(\cdot|u,s,h,r_{t})$ /* Builder writes code $x_{t}$/
       $(r_{t},v_{t})\sim\mathrm{I}(\cdot|s,x)$ /* Inspector checks code, outputs verdict $v_{t}$ and suggestion $r_{t}$/
       $t\leftarrow t+1$;
end while
return x
Algorithm 2 Generate Code With Inspection
@@ 3.4. Compilation, Save and Reload
Skill Library Workflow. This image depicts the workflow of a ’Skill Library’ module in a software context. The diagram is split into two main sections, ’Input’ on the left and ’Output’ on the right, with an arrow labeled ’GPT-4’ pointing from the left section to the right, indicating the flow of processing. The ’Input’ section shows a simplified user interface with buttons for ’Scene,’ ’SA,’ ’User,’ ’P,’ ’SL,’ ’Developer,’ and ’Server,’ and below is a text box with the user prompt ’create a whale and make it swim happily.’ Below this, a list of skills is detailed, including ’ObjectRetriever,’ ’Animation,’ ’GetMeshSize,’ and ’MRTK,’ each with a brief description of its function. The ’Output’ section contains code snippets and references to ’Object Retriever’ and ’Animation,’ suggesting these are the selected outputs from the Skill Library module. The bottom of both sections restates the user prompt and mentions that the relevant scene and object are not applicable.
@@ 3.5. Skill Library
The creation of the Skill Library Module is motivated by two primary challenges. The first is the token size limitation imposed by the GPT architecture on the context, or the ”metaprompt,” provided to the Builder. Typically, the Builder is presented with a comprehensive list of various APIs and plug-ins that could be employed to meet the user’s needs. As the range of available skills expands, this list lengthens, eventually surpassing GPT’s token size limit for public users. The second challenge lies in the Builder’s attention capacity, which appears to be limited. Even when we attempt to condense all the available skills into the Builder’s metaprompt, it struggles to keep track of a specific skill when the list becomes too lengthy. This limitation is further exacerbated by the necessity to include precise coding examples for each plugin to ensure their effective utilization by GPT. To address these challenges, we created the Skill Library module, denoted as $\mathrm{L}(h|u)$, which serves as a centralized repository for all available skills and as an attention mechanism that retrieves only the skills relevant to a specific user prompt. We illustrate this module in Figure 6.
@@
As an illustrative example, consider a skill we created for GPT’s use, which leverages a combination of generative and contrastive models along with the Sketchfab API to source and integrate 3D models into a scene. We have also created skills that allow the generation of animation of a rigged object in real-time (Huang et al., 2023a). While we delve into the specifics of a skill in the next section, it is worth noting that the Skill Library only receives a high-level summary of how this particular skill functions, along with similar descriptors for other skills. The actual examples needed to use this skill are then retrieved and supplied to the Builder for execution.
@@
This approach ensures that the Skill Library and the Builder work in tandem to efficiently and effectively generate code that fulfills the user’s request while overcoming the token size and attention capacity limitations of LLMs.
Object Retriever pipeline for generating a 3D scene. This diagram depicts the Object Retriever pipeline for generating a 3D scene based on a user prompt. The input stage begins with a textual prompt for a scene with specific objects. For each object, such as a clock, DALL-E 2 generates a target 3D image. Simultaneously, the Sketchfab API retrieves multiple potential matches for each object, displayed as thumbnails. CLIP then processes these images to produce embeddings that include the target image. The top 5 candidates are selected based on text similarity, and the final choice is made based on the highest visual similarity to the target image. This process is repeated for each object in the prompt. The final output is a rendered 3D scene on the right, featuring a wall clock, a picture frame on a wall, a wooden chair, and a table with an apple on it, all placed within a room.
@@ 4. Incorporating Existing Open-Source 3D Assets
The process of generating interactive 3D scenes often involves the creation and placement of various objects. For instance, a request to create an office space might be decomposed into the generation of a desk, a chair, a lamp, and a clock. While it is possible to generate these objects using primitives, a method that works well even for composite objects like a car or an entire room (depicted in the car of Figure 8 and the kitchen of Figure 1), there is a need to leverage the intricate objects created by artists and 3D developers that exhibit high real-world fidelity. Previous work has utilized objects from Sketchfab (Roberts et al., 2022b, a) and used the priors of GPT to size them accordingly to the real world. However, this approach encounters challenges when the user prompts an object, say a clock, and Sketchfab offers 50 different clocks, only three of which are suitable for an office setting.
@@
This process is repeated to generate entire scenes. Algorithm 3 and Figure 7 describe this pipeline. There is potential for further exploration to improve this pipeline. For instance, selecting from the visual similarity space before the language similarity space might yield better results. Future work will involve human feedback to identify the workflow that maximizes the likeness between the 3D object loaded and the user’s intended object.
Input : $u$: user’s prompt
Require : $o$: object in $u$;
$T$: target image;
$\bm{S}$: screenshots;
$L(o,s_{i})$: language similarity;
$V(T,s_{i})$: visual similarity.
$\bm{S}^{\prime}\leftarrow\text{Top 5}s_{i}\in\bm{S}\text{with highest}L(o,s_{i})$;
$s^{*}\leftarrow\arg\max_{s_{i}\in\bm{S^{\prime}}}V(T,s_{i})$;
return $s^{*}$
@@
Cross-Platform and Cross-Scene Transferability made possible by LLMR. This composite image illustrates the transferability of a car model across different platforms and scenes using a specific framework. The first panel on the left, labeled ’A’ under the VR category, shows a simple car model created using Unity primitives with a red body, wheels, black, and headlights in yellow, indicating it can be controlled with keyboard inputs. The middle panel, labeled ’B’, under the VR category, depicts the same car model placed on a lunar terrain with moon-like gravity within Unity, demonstrating the model’s adaptability to different virtual environments. The rightmost panel, labeled ’C’ under the Mobile + AR category, showcases the car in a real-world setting overlayed through augmented reality (AR) on a mobile device, showing the car’s ability to collide with physical objects and be controlled using the phone’s Inertial Measurement Unit (IMU) data.
Sketching objects into existence with LLMR. Three-part image demonstrating the LLMR process for creating virtual objects. Panel A shows a virtual reality (VR) environment with a simple, white, floating paintbrush next to a VR controller, illustrating the user’s request for a ”magic paintbrush” to be attached to the controller. Panel B depicts the user’s action within the VR environment, showcasing a complex, scribbled blue drawing in three-dimensional space, representing the automatic conversion of a line renderer into a functioning paintbrush. Panel C presents a variety of chair models in a 3D space, indicating the transformation from a 2D sketch to 3D objects using 2D-3D ControlNet and an API, which allows for the generation and interaction with multiple chair models in different VR platforms.
@@ 5. Memory Management
By default, language models generate new words based on all previously sampled tokens, a configuration that may not be ideal due to their finite context length. For instance, this may hinder the model’s ability to engage in extended conversations. To mitigate this, techniques such as dialogue summarization and distillation can be employed (Wang et al., 2023a; Bae et al., 2022; Huang et al., 2023b). Additional research has delved into leveraging persistent memory and retrieving in-context examples from databases to enhance few-shot performance (Wang et al., 2023b; Zhong et al., 2023).
Module Memory Mode
@@
Memory-less
Scene Analyzer
Memory-less
@@
Figure 10. Accessible Interface Features in Action. A1 and A2 show how a user can prompt the system to adjust the color scheme of a kitchen scene for red-green color-blind compatibility. B1 and B2 demonstrate the activation of a magnifier tool. C1 and C2 reveal the option to hide objects deemed not kid-friendly.
Accessible Interface Features in Action. Six-panel image sequence illustrating accessible interface features within a kitchen scene. Panels A1 and A2 display the same corner of a modern kitchen, with the color scheme adjusted from standard to a palette suitable for red-green color-blind individuals, showing a marked difference in contrast and color tones. Panels B1 and B2 highlight a digital magnifying tool in use; B1 shows a standard view of the kitchen counter and stove, while B2 zooms in on the stove controls for clearer visibility. Panels C1 and C2 focus on child safety features; C1 depicts the kitchen with various items on the counter, including knives, while C2 shows the same view with potentially dangerous objects like knives hidden from sight, creating a safer environment for children.
Figure 10. Accessible Interface Features in Action. A1 and A2 show how a user can prompt the system to adjust the color scheme of a kitchen scene for red-green color-blind compatibility. B1 and B2 demonstrate the activation of a magnifier tool. C1 and C2 reveal the option to hide objects deemed not kid-friendly.
@@ 6.1. Installation
Our framework can be easily added to any existing Unity scenes. The framework consists of a unity package and a few additional open-sourced packages (such as GLTF loader and OpenAI), and the installation process takes only a few steps. This enables anyone with an OpenAI API key to try our framework. We are strong proponents of the adaptability of our framework, and so we have open-sourced the foundational framework along with several examples on GitHub (https://llm4mr.github.io/). Readers who wish to try our framework can try out the example playground scenes or can easily add our Unity package to their existing Unity projects. They would need to obtain an OpenAI API key, a copy of the Roslyn compiler and optionally an account for Sketchfab if they wish to automatically load existing assets. In the Appendix, we also provide the metaprompts used for each LLMR’s modules for transparency.
Figure 11. Spontaneous Creation of Teaching Guides. A demonstration of creating a guide for operating a coffee machine in which LLMR animates a hand model to point out the various steps of the operation. Our framework allows for the rapid creation of such guides and furthermore allows users to ask questions that were not predicted by the instructor beforehand, with appropriate motions being animated on the fly.
Spontaneous Creation of Teaching Guides. Four-panel image sequence demonstrating an interactive guide for operating a coffee machine with LLMR technology. The top two panels, labeled A1 and A2, represent the instructor’s view. A1 shows a coffee machine interface with a virtual hand model hovering over a button, indicating the first step with the text prompt ”Load the coffee machine and a hand model.” A2 depicts the hand model gesturing to a steamer knob with the text ”Implement the following steps, using hand gestures to visualize…” The bottom two panels, labeled B1 and B2, depict the trainee’s view with a timer and query responses. B1 shows the hand model pressing a button with a timer displaying 42:53 and the caption ”OK, How long should I press the button?” B2 illustrates the hand model turning the steamer knob with the question ”Which direction should I rotate the steamer?” This visual aid conveys the dynamic nature of the teaching guide, with a system that animates hand motions in response to user queries, facilitating on-the-fly instructional interactions.
Figure 11. Spontaneous Creation of Teaching Guides. A demonstration of creating a guide for operating a coffee machine in which LLMR animates a hand model to point out the various steps of the operation. Our framework allows for the rapid creation of such guides and furthermore allows users to ask questions that were not predicted by the instructor beforehand, with appropriate motions being animated on the fly.
@@ 7.1. Game Design and Creativity
Figure 12. Simulated Rescue Plan. The HoloLens 2 displays the automated generation of a simulation of a rescue plan using our framework. The guide shows an interactable 3D terrain, helicopter, and simulated wind, allowing rescue workers to visualize the flight path under different weather conditions.
Simulated Rescue Plan. The image shows a side-by-side comparison labeled as (A) and (B). Both depict a simulated rescue plan using augmented reality on a HoloLens device. Description for (A): A virtual 3D environment is overlaid on a physical office setting. Four small, red spheres are arranged in a square pattern on the floor. A virtual helicopter is positioned near the center, slightly above a black office chair. The text instruction suggests interaction with the terrain and helicopter, and the presence of a wind effect represented by a red arrow near the bottom left sphere. Description for (B): Similar to (A), with the virtual helicopter now located closer to the rightmost red sphere. The red arrow indicating the wind effect is more pronounced and closer to the sphere, which could suggest the helicopter has moved due to the simulated wind.
Figure 12. Simulated Rescue Plan. The HoloLens 2 displays the automated generation of a simulation of a rescue plan using our framework. The guide shows an interactable 3D terrain, helicopter, and simulated wind, allowing rescue workers to visualize the flight path under different weather conditions.
@@ 7.2. Accessibility and Adaptive Interface
Similar to the accessibility feature in 2D documents, our framework can also be prompted to make a 3D scene accessible and adaptive to different user needs and preferences. Figure 10 shows three examples of editing an existing virtual kitchen scene to different requests. For example, one can request to make the scene to be more friendly to red-green color-blind users. For someone who is near-sighted, they can prompt a magnifier tool that zooms into a particular part of the room. An architect can use our framework to figure out if the space is friendly for wheelchair users or make sure objects in the room are child-proof. These examples show how our framework leverages LLMs’ prior knowledge and puts the knowledge into the context of a spatial world at a human scale.
@@ 7.3. Remote Assistance and Planning
In a remote training scenario, typically, creating such a 3D interactive training guide requires custom creation, from rigging a gesture to placing a UI element. An instructor can use our framework to automate the generation of a training guide from a list of instructions. (Figure 11). The trainee can then, for example, use an AR device that overlays information on the machine. As the trainee advances through the steps, they can ask questions directly to the guide where answers can be generated in the context of the trainee’s learning progress. In another scenario of remote rescue planning, helicopter operators can prompt a simulation of the flight path given several target locations and see how the flight path might be affected by different wind conditions (Figure 12).
Figure 13. An illustration of our experimental setup. We provide the bathroom scene (left) and a subset of the 150 prompts (right) used in this space for the evaluation provided in Figure 14.
An illustration of our ablation experimental setup. The image is split into two sections: ”Existing Scene” on the left and ”Example Prompts” on the right. The photo displays a modern bathroom scene with a neutral color palette. A freestanding white bathtub is placed in the center, with a grey-tiled wall behind it featuring three recessed shelves adorned with plants and towels. To the left of the bathtub, a towel radiator is mounted on the wall, and to the right, a large mirror hangs above a sink with visible faucets. The floor is also grey, complementing the wall tiles. On the right side, a list of example interactive prompts is displayed against a white background. The prompts include actions such as creating a sphere on top of the bathtub, changing the toilet’s color to gold, and letting water flow out of the sink faucet when clicked. Other prompts suggest more complex interactions like filling the bathtub with water and making the room more relaxing, indicating a dynamic and interactive virtual environment.
Figure 13. An illustration of our experimental setup. We provide the bathroom scene (left) and a subset of the 150 prompts (right) used in this space for the evaluation provided in Figure 14.